OpenStudy (anonymous):

square root 8a^8 ....how do I do this by hand??

OpenStudy (anonymous):

The answer is \[2\sqrt{2a ^{4}}\]

OpenStudy (anonymous):

like this? \[\sqrt{8a^8}\]

OpenStudy (anonymous):

but I need to know how to do the steps...Yes Joemath like that!! :)

OpenStudy (anonymous):

to be able to pull things out of a square root, you need them to be squares. For example: \[\sqrt{36} = \sqrt{6^2} = 6\] or \[\sqrt{50} = \sqrt{25*2} = \sqrt{5^2*2} = 5\sqrt{2}\] So lets see what squares we can find in your problem: \[\sqrt{8a^8} = \sqrt{4*2(a^4)^2} =\sqrt{(2^2)*2*(a^4)^2}\] the squares we can pull out are a 2, and a^4, so we pull them out, and leave everything else behind: \[\sqrt{2^2*2*(a^4)^2} = 2a^4\sqrt{2}\] that would be the final answer.

OpenStudy (anonymous):

Thank you, that made sense :)!!